OpenStudy (anonymous):

What is a polynomial function in standard form with the zeros 1, 2, -2, and -3?

OpenStudy (anonymous):

I know we start by doing (x-1)(x-2)(x+2)(x+3)

OpenStudy (anonymous):

the one that looks like \[\large (x-1)(x-2)(x+2)(x+3)\] after you multiply out

OpenStudy (anonymous):

i would start with \[(x-2)(x+2)=x^2-4\]

OpenStudy (anonymous):

I don't know how to multiply it out... That is where I get confused.

OpenStudy (anonymous):

then \[(x-1)(x+3)=x^2+3x-x-3=x^2+2x-3\] and finally multiply \[(x^2-4)(x^2+2x-3)\]

OpenStudy (anonymous):

course you can do them in any order you like, but that is what i think would be easiest want to see what is really easiest??

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Ohhh, I get it now! I got as far as x^4+2x^3, but then I got confused but I get it now. Thank you!